# ReSem3D: Refinable 3D Spatial Constraints via Fine-Grained Semantic Grounding for Generalizable Robotic Manipulation

### v2

@@ Abstract
Note to Practitioners—This work is motivated by the practical challenge of enabling robots to perform robust and generalizable manipulation in semantically diverse environments such as households and chemical laboratories. Existing systems frequently fall short in fine-grained semantic modeling of spatial constraints and real-time adaptability, limiting their effectiveness in real-world applications. We propose a unified framework that leverages recent advances in multimodal language and vision models to automatically construct 3D spatial constraints from natural language instructions and visual observations. These constraints are hierarchically refined and integrated into a real-time control system that supports reactive and context-aware behavior. The proposed method is then validated in both simulated and physical environments, demonstrating strong generalization across object categories and task scenarios, which enables robotic systems to perform complex, zero-shot tasks in dynamic and semantically diverse environments, overcoming previous limitations in coarse constraint modeling. This framework offers a practical solution for implementing adaptable manipulation systems without cumbersome task-specific manual programming.
Index Terms:
Multimodal Large Language Models, Vision Foundation Models, Task and Motion Planning, Manipulation.
@@ I Introduction
With the advancement of robotic autonomy in unstructured environments, the growing diversity of semantic descriptions and the complexity of visual modeling have become critical challenges limiting system generalization and robustness [1]. Spatial constraints [2, 3] serve as the critical interface between high-level semantic representations and low-level action spaces, enabling the mapping of abstract semantics into executable actions. Recent advances in cross-modal reasoning powered by Multimodal Large Language Models (MLLMs) [4] and Vision Foundation Models (VFMs) [5, 6] have significantly improved visual grounding capability, which is crucial for constructing accurate 3D spatial constraints. For instance, tweezers-based grasping entails localizing the coarse semantic grasp region and reasoning about the 3D geometric relationships between the tweezer tips and the grasp point. Based on these inferred relations, spatial constraint models can be constructed and parsed in real-time by the constraint optimizer, thereby enabling closed-loop control and fine-grained manipulation in dynamic environments.
@@
However, current methods lack an optimizable intermediate representation that effectively links linguistic intent with the action space, thereby hindering consistent alignment between semantic goals and physical execution. As a core intermediate representation, 3D spatial constraints parameterize the mapping from semantic intent to executable actions through a set of task-grounded poses. Nevertheless, the geometric reasoning and spatial localization capabilities of MLLMs remain insufficient for directly deriving fine-grained constraints from language. Prior efforts have leveraged MLLMs and VFMs to enhance language–vision semantic alignment [7, 8], facilitating the construction of 3D spatial constraints for motion planning [9, 10]. These constraints often remain limited to coarse localization and lack semantic-guided refinement. Furthermore, downstream controllers typically rely on task-space planning, which is limited by the stability and efficiency of inverse kinematics solvers. These limitations raise two critical challenges: 1) how to design a flexible and optimizable 3D spatial constraint model that automatically maps semantic intent to robotic actions, and 2) how to ensure stable closed-loop and real-time execution within the low-level action space in semantically diverse environments.
@@
To overcome the first challenge, we propose a two-stage hierarchical 3D spatial constraint modeling approach driven by the synergistic reasoning of VFMs and MLLMs: part-level constraint extraction and region-level constraint refinement. The part-level extraction leverages VFMs to identify semantically relevant part-level regions from the RGB observation, while generating explicit digital markers as visual prompts to facilitate MLLMs in coarse-grained constraint extraction. The region-level refinement builds dense semantic grids within masked regions, enabling fine-grained semantic grounding of geometric features by MLLMs. Subsequently, 3D spatial constraints are encoded as cost functions parsed in real-time, and supplied to downstream optimizers. To tackle the second challenge, inspired by a finite-horizon optimal control approach grounded in a general-purpose physics simulator [11, 12], we implement a real-time Model Predictive Path Integral (MPPI) control algorithm in the Isaac Gym platform [13], which delivers 15 Hz joint-space velocity commands instead of task-space commands for stable closed-loop execution.
@@
In addition, we develop a Task and Motion Planning (TAMP) framework that extends the hierarchical recursive prompting structure of Liang et al. [14], driven by MLLM-based reasoning, enabling autonomous and closed-loop manipulation from perception to action. The framework demonstrates semantics-driven adaptability by autonomously determining the required granularity of 3D spatial constraints from natural language instructions. It decomposes multi-stage tasks and generates precondition checks, postcondition evaluations, and optimizable cost functions necessary for the operation of each subtask. Building on this, a subtask-level backtracking strategy is formulated, leveraging precondition feasibility to initiate cross-stage replan, enabling behavior recovery and reactive control under dynamic perturbations.
@@ II-A Visual Grounding with Foundation Models
Visual grounding refers to mapping semantic features from natural language to image regions and spatial entities, evolving from low-level perception to semantic spatial reasoning with the advancement of foundation models. Prior VFMs focused on robust image perception, leveraging self-supervised learning [15, 16] and architectures including CNNs [17, 18] and Vision Transformers [19], which produce visual encodings for downstream tasks like object detection [20, 21], segmentation [22, 23], and feature extraction [24]. Furthermore, pixel-level parsing was enhanced through latent feature encoding and embedded visual prompts [25, 26]. Subsequently, Vision-Language Models (VLMs) perform multimodal semantic grounding through visual-linguistic representations, achieving notable progress in image-text retrieval [27] and open-vocabulary object detection [28], but still struggle with fine-grained visual grounding. Building on this, researchers introduced Large Language Models (LLMs) [29] into vision-language systems, leading to MLLMs [30] that incorporate visual features as context for multimodal reasoning. Despite this, MLLMs remain dependent on front-end visual encoders, limiting spatial precision and semantic understanding. Recent studies have explored the synergy between VFMs and MLLMs by embedding visual prompts for semantic alignment [7], but fine-grained visual grounding remains a critical limitation. To address this challenge, we propose a multi-granularity semantic understanding framework with hierarchical multimodal reasoning, featuring part-level extraction and region-level refinement to enable fine-grained visual grounding.
@@ II-B Spatial Constraints in Robotic Manipulation
Spatial constraints are essential for robotic manipulation, bridging high-level semantics and low-level actions to ensure task feasibility. Conventional methods construct constraints through explicit geometric modeling of object shapes, contact dynamics, and environmental structure [31, 32] to enable motion planning. While physically interpretable and effective in structured scenarios, their dependence on accurate modeling constrains their flexibility in complex settings. Recently, perception-driven learning has been used to infer potential manipulation regions from visual observations, including affordance recognition [33], grasp pose detection [34, 35], target pose estimation [36], and semantic keypoint generation [37]. These approaches have further extended to multimodal manipulation frameworks that incorporate foundation model reasoning [38, 39]. While capable of adapting to object deformations and semantic changes, they still require manual data collection and struggle with fine-grained semantic parsing and precise modeling of complex geometric relationships. Current approaches integrate language and vision models to enable zero-shot semantic region perception [2, 9]. Nevertheless, aforementioned methods typically lack fine-grained semantic localization and executable geometric specifications. To overcome these limitations, we propose a two-stage 3D spatial constraint modeling framework that hierarchically maps natural language instructions to semantic objectives, enabling fine-grained and executable spatial constraints.
@@ II-C Task and Motion Planning
Task and Motion Planning (TAMP) provides a pivotal framework connecting high-level task reasoning with low-level motion execution. Conventional approaches often leverage formal languages such as PDDL and HTN to model symbolic tasks, solving task sequences and motion trajectories through logic-geometric and mixed-integer programming [40, 41]. Although offering strong interpretability, these methods rely on manually designed task models and action primitives, limiting their adaptability in open environments. With the advancement of LLMs, recent works have explored zero-shot task planning based on predefined motion primitives [42]. However, these approaches lack effective modeling of geometric constraints and environmental dynamics. Building on these developments, recent research has explored integrating LLMs with motion planning to enable high-level task decomposition and continuous action generation without predefined primitives [10]. Although such methods enhance the flexibility of TAMP, they often depend on online optimization in task space, posing challenges for real-time execution. In order to address these challenges, we employ a real-time MPPI optimizer built on Isaac Gym, leveraging joint velocity control to enhance reactive behavior during execution. Concurrently, subtask-level constraints and optimization objectives are automatically generated by the MLLM, enabling effective closed-loop control and zero-shot adaptation in dynamic environments.
@@ III-A ReSem3D Problem Formulation
As shown in Fig. 1, given a free-form natural language instruction =$\mathcal{L}=\{L_{1},L_{2},\ldots,L_{n}\}$ from the user, $\mathcal{L}$ typically exhibits significant long-term dependencies and complex semantic reasoning. Directly generating complete 3D spatial constraints and optimizing them both compromises reasoning accuracy and complicates constraint modeling. Therefore, ReSem3D employs a hierarchical modeling strategy by decomposing the task into sequential subtasks and performing constraint reasoning and optimization progressively for each $L_{i}$, $i\in\{1,2,\ldots,n\}$.
@@ III-B1 Part-Level Constraint Extraction
[…]
@@
Step 2: Part-Level Semantic Consistency Clustering. Density-based clustering (DBSCAN) [43] is performed on the filtered mask set to further reduce visual ambiguity and redundancy, with masks of identical semantic labels merged within each cluster. The clustered mask set is denoted as
@@ III-B2 Region-Level Constraint Refinement
Step 4: Fine-grained semantic reasoning by MLLM. Achieving region-level constraint refinement and deriving new labels ^${\hat{\ell}}_{k,s}^{j}$ based on the target geometric involve feeding the original RGB image $I_{\text{rgb}}$ , the original mask ${\widetilde{m}}_{j}$ , and the target mask ${\widetilde{m}}_{j}$ as visual prompts into the MLLM. Subsequently, the numerical label $z_{j}$ and the semantic label $\ell_{k,s}^{j}$ are presented as textual prompts to the MLLM for reasoning.
Step 5: Region-Level Constraint Extraction. The refined constaint is constructed by inversely mapping the labels ^${\hat{\ell}}_{k,s}^{j}$ to their corresponding grids $\mathcal{G}_{k,s}^{j}$ via the instantiated mapping $\mathcal{U}:\mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$. The associated grid centroids $\bm{c}_{j}^{k,s}$ are then projected to the original mask coordinate as $\bm{c}_{j}$ using Equations (17) and (18). Finally, the 2D coordinates are transformed into 3D space under the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, thereby forming the initial constraint $\mathcal{C}_{i}^{\text{init}}$.
@@
[…]
@@ III-C Real-Time Close-Loop Control Strategy
To solve the constrained optimization problem in Equation (2) with real-time performance, the Model Predictive Path Integral (MPPI) control algorithm is employed on the Isaac Gym simulation platform. MPPI is a stochastic optimization method designed for robot control in dynamic and uncertain environments. Essentially, the algorithm samples multiple candidate control sequences from the control input space, evaluates the cost functions based on simulated state trajectories, and updates the control policy through importance-weighted averaging. The GPU-accelerated parallel simulation enabled by Isaac Gym significantly enhances sampling efficiency, thereby supporting real-time execution in dynamic environments. For a detailed explanation of MPPI, please refer to [44, 45]. In this work, we derive the specific formulation for solving the constrained optimization problem presented in Equation (2).
@@
where $\lambda_{p}$ and $\lambda_{r}$ are the weights for position and orientation costs, $\lambda_{c}$ denote the weight for the collision cost $\mathcal{D}_{c}$, which is formulated based on the ESDF representation, as described in [9]. In this work, the objective function is subject to preconditions and postconditions as specified in Equation (2), the precondition typically ensures that the end-effector is within the designated region prior to the execution stage, while the postcondition assesses whether the end-effector has successfully reached the desired pose, defined as
@@ III-D MLLM-driven Automated Modeling of TAMP
As depicted in Fig. 4, the Language Model Program (LMP) framework is built on a hierarchical recursive architecture [14], enabling structured invocation and multi-level prompt nesting. Based on the framework, we have designed and encapsulated seven LMP modules within ReSem3D, corresponding respectively to Task Planner, Constraint Extraction, Constraint Refinement, Subtask Execution, Preconditions Building, Cost Function Building, and PostConditions Building. Each module implements an independent prompt structure and interfaces with system components through standardized APIs, enabling the automated modeling for task and motion planning.
@@ IV-A Experimental Setup
Hardware Configuration. As shown in Fig. 6, our real-world experimental platform is constructed around the 6-DOF Universal Robots UR5e collaborative robotic arm, equipped with a PGI-140-80 two-finger parallel gripper and an Intel RealSense D435i RGB-D camera. For simulation, we employ the Omnigibson [46] platform developed by Stanford University, and utilize both the UR5e and the 7-DOF Franka Emika Panda collaborative arms.
@@
Baselines. We compare against three representative frameworks: VoxPoser [10], ReKep [9] and CoPa [2]. VoxPoser integrates large language and vision models to generate 3D value maps interpretable by robots, enabling closed-loop trajectory synthesis without predefined motion primitives. It leverages GPT-4 and a visual frontend for semantic grounding of task objects, supporting flexible manipulation across diverse tasks. ReKep fuses VLMs and MLLMs to infer semantic keypoint constraints for structured tasks and solves end-effector trajectories in real time via hierarchical optimization. CoPa leverages foundation vision-language models to propose a two-stage task execution process, performing task-oriented grasping via coarse-to-fine visual grounding in the first stage, and identifying task-relevant spatial and geometric constraints for robot planning in the second stage. All three frameworks deeply integrate large language and vision models, employing semantics-driven visual grounding and control optimization to achieve efficient and robust robotic manipulation.
@@
VFM and MLLM. We utilize GPT-4o [4] from OpenAI as a callable MLLM and combine it with the visual foundation model FastSAM [5] to enable high-efficiency multimodal reasoning (prompt design is available at ReSem3D.github.io. During the dynamic disturbance phase, we incorporate the TAPNext model from Google DeepMind [47] together with real-time RGB frame input to achieve online point tracking at approximately 20 Hz. This facilitates real-time inference of evolving spatial constraints in dynamic environments, thereby enhancing the stability and robustness of task execution.

### v3

@@ I Introduction
With the advancement of robotic autonomy in unstructured environments, the growing diversity of semantic descriptions and the complexity of visual modeling have become critical challenges limiting system generalization and robustness [ref1]. Spatial constraints [ref2, ref3] serve as the critical interface between high-level semantic representations and low-level action spaces, enabling the mapping of abstract semantics into executable actions. Recent advances in cross-modal reasoning powered by Multimodal Large Language Models (MLLMs) [ref4] and Vision Foundation Models (VFMs) [ref5, ref6] have significantly improved visual grounding capability, which is crucial for constructing accurate 3D spatial constraints. For instance, tweezers-based grasping entails localizing the coarse semantic grasp region and reasoning about the 3D geometric relationships between the tweezer tips and the grasp point. Based on these inferred relations, spatial constraint models can be constructed and parsed in real-time by the constraint optimizer, thereby enabling closed-loop control and fine-grained manipulation in dynamic environments.
@@
However, current methods lack an optimizable intermediate representation that effectively links linguistic intent with the action space, thereby hindering consistent alignment between semantic goals and physical execution. As a core intermediate representation, 3D spatial constraints parameterize the mapping from semantic intent to executable actions through a set of task-grounded poses. Nevertheless, the geometric reasoning and spatial localization capabilities of MLLMs remain insufficient for directly deriving fine-grained constraints from language. Prior efforts have leveraged MLLMs and VFMs to enhance language–vision semantic alignment [ref7, ref8], facilitating the construction of 3D spatial constraints for motion planning [ref9, ref10]. These constraints often remain limited to coarse localization and lack semantic-guided refinement. Furthermore, downstream controllers typically rely on task-space planning, which is limited by the stability and efficiency of inverse kinematics solvers. These limitations raise two critical challenges: 1) how to design a flexible and optimizable 3D spatial constraint model that automatically maps semantic intent to robotic actions, and 2) how to ensure stable closed-loop and real-time execution within the low-level action space in semantically diverse environments.
@@
To overcome the first challenge, we propose a two-stage hierarchical 3D spatial constraint modeling approach driven by the synergistic reasoning of VFMs and MLLMs: part-level constraint extraction and region-level constraint refinement. The part-level extraction leverages VFMs to identify semantically relevant part-level regions from the RGB observation, while generating explicit digital markers as visual prompts to facilitate MLLMs in coarse-grained constraint extraction. The region-level refinement builds dense semantic grids within masked regions, enabling fine-grained semantic grounding of geometric features by MLLMs. Subsequently, 3D spatial constraints are encoded as cost functions parsed in real-time, and supplied to downstream optimizers. To tackle the second challenge, inspired by a finite-horizon optimal control approach grounded in a general-purpose physics simulator [ref11, ref12], we implement a real-time Model Predictive Path Integral (MPPI) control algorithm in the Isaac Gym platform [ref13], which delivers 15 Hz joint-space velocity commands instead of task-space commands for stable closed-loop execution.
@@
In addition, we develop a Task and Motion Planning (TAMP) framework that extends the hierarchical recursive prompting structure of Liang et al. [ref14], driven by MLLM-based reasoning, enabling autonomous and closed-loop manipulation from perception to action. The framework demonstrates semantics-driven adaptability by autonomously determining the required granularity of 3D spatial constraints from natural language instructions. It decomposes multi-stage tasks and generates precondition checks, postcondition evaluations, and optimizable cost functions necessary for the operation of each subtask. Building on this, a subtask-level backtracking strategy is formulated, leveraging precondition feasibility to initiate cross-stage replan, enabling behavior recovery and reactive control under dynamic perturbations.
@@ II-A Visual Grounding with Foundation Models
Visual grounding refers to mapping semantic features from natural language to image regions and spatial entities, evolving from low-level perception to semantic spatial reasoning with the advancement of foundation models. Prior VFMs focused on robust image perception, leveraging self-supervised learning [ref15, ref16] and architectures including CNNs [ref17, ref18] and Vision Transformers [ref19], which produce visual encodings for downstream tasks like object detection [ref20, ref21], segmentation [ref22, ref23], and feature extraction [ref24]. Furthermore, pixel-level parsing was enhanced through latent feature encoding and embedded visual prompts [ref25, ref26]. Subsequently, Vision-Language Models (VLMs) perform multimodal semantic grounding through visual-linguistic representations, achieving notable progress in image-text retrieval [ref27] and open-vocabulary object detection [ref28], but still struggle with fine-grained visual grounding. Building on this, researchers introduced Large Language Models (LLMs) [ref29] into vision-language systems, leading to MLLMs [ref30] that incorporate visual features as context for multimodal reasoning. Despite this, MLLMs remain dependent on front-end visual encoders, limiting spatial precision and semantic understanding. Recent studies have explored the synergy between VFMs and MLLMs by embedding visual prompts for semantic alignment [ref7], but fine-grained visual grounding remains a critical limitation. To address this challenge, we propose a multi-granularity semantic understanding framework with hierarchical multimodal reasoning, featuring part-level extraction and region-level refinement to enable fine-grained visual grounding.
@@ II-B Spatial Constraints in Robotic Manipulation
Spatial constraints are essential for robotic manipulation, bridging high-level semantics and low-level actions to ensure task feasibility. Conventional methods construct constraints through explicit geometric modeling of object shapes, contact dynamics, and environmental structure [ref31, ref32] to enable motion planning. While physically interpretable and effective in structured scenarios, their dependence on accurate modeling constrains their flexibility in complex settings. Recently, perception-driven learning has been used to infer potential manipulation regions from visual observations, including affordance recognition [ref33], grasp pose detection [ref34, ref35], target pose estimation [ref36], and semantic keypoint generation [ref37]. These approaches have further extended to multimodal manipulation frameworks that incorporate foundation model reasoning [ref38, ref39]. While capable of adapting to object deformations and semantic changes, they still require manual data collection and struggle with fine-grained semantic parsing and precise modeling of complex geometric relationships. Current approaches integrate language and vision models to enable zero-shot semantic region perception [ref2, ref9]. Nevertheless, aforementioned methods typically lack fine-grained semantic localization and executable geometric specifications. To overcome these limitations, we propose a two-stage 3D spatial constraint modeling framework that hierarchically maps natural language instructions to semantic objectives, enabling fine-grained and executable spatial constraints.
@@ II-C Task and Motion Planning
Task and Motion Planning (TAMP) provides a pivotal framework connecting high-level task reasoning with low-level motion execution. Conventional approaches often leverage formal languages such as PDDL and HTN to model symbolic tasks, solving task sequences and motion trajectories through logic-geometric and mixed-integer programming [ref40, ref41]. Although offering strong interpretability, these methods rely on manually designed task models and action primitives, limiting their adaptability in open environments. With the advancement of LLMs, recent works have explored zero-shot task planning based on predefined motion primitives [ref42]. However, these approaches lack effective modeling of geometric constraints and environmental dynamics. Building on these developments, recent research has explored integrating LLMs with motion planning to enable high-level task decomposition and continuous action generation without predefined primitives [ref10]. Although such methods enhance the flexibility of TAMP, they often depend on online optimization in task space, posing challenges for real-time execution. In order to address these challenges, we employ a real-time MPPI optimizer built on Isaac Gym, leveraging joint velocity control to enhance reactive behavior during execution. Concurrently, subtask-level constraints and optimization objectives are automatically generated by the MLLM, enabling effective closed-loop control and zero-shot adaptation in dynamic environments.
@@ III-A ReSem3D Problem Formulation
As shown in Fig. 1, given a free-form natural language instruction $\mathcal{L}=\{L_{1},L_{2},\ldots,L_{n}\}$ from the user, $\mathcal{L}$ typically exhibits significant long-term dependencies and complex semantic reasoning. Directly generating complete 3D spatial constraints and optimizing them both compromises reasoning accuracy and complicates constraint modeling. Therefore, ReSem3D employs a hierarchical modeling strategy by decomposing the task into sequential subtasks and performing constraint reasoning and optimization progressively for each $L_{i}$, $i\in\{1,2,\ldots,n\}$.
@@ III-B1 Part-Level Constraint Extraction
As illustrated in Fig. 2, given subtask instruction $L_{i}$ and RGB-D observation $O_{i}\in\mathbb{R}^{H\times W\times 4}$, we apply the pretrained image segmentation model [ref5] to produce the initial set of masks $M=\{m_{1},m_{2},\ldots,m_{g}\}$, where $m_{j}\in{\{0,\ 1\}}^{H\times W}$ and $\forall j\in\{1,\ldots,g\}$. Here, $H$ and $W$ denote the height and width of the image, respectively. To extract high-quality and discriminative part-level constraints, we design the following three-step post-processing strategy.
@@
Step 2: Part-Level Semantic Consistency Clustering. Density-based clustering (DBSCAN) [ref43] is performed on the filtered mask set to further reduce visual ambiguity and redundancy, with masks of identical semantic labels merged within each cluster. The clustered mask set is denoted as
@@ III-B2 Region-Level Constraint Refinement
Step 4: Fine-grained semantic reasoning by MLLM. Achieving region-level constraint refinement and deriving new labels ${\hat{\ell}}_{k,s}^{j}$ based on the target geometric involve feeding the original RGB image $I_{\text{rgb}}$ , the original mask ${\widetilde{m}}_{j}$ , and the target mask ${\widetilde{m}}_{j}$ as visual prompts into the MLLM. Subsequently, the numerical label $z_{j}$ and the semantic label $\ell_{k,s}^{j}$ are presented as textual prompts to the MLLM for reasoning.
Step 5: Region-Level Constraint Extraction. The refined constaint is constructed by inversely mapping the labels ${\hat{\ell}}_{k,s}^{j}$ to their corresponding grids $\mathcal{G}_{k,s}^{j}$ via the instantiated mapping $\mathcal{U}:\mathcal{G}_{k,s}^{j}\rightarrow\ell_{k,s}^{j}$. The associated grid centroids $\bm{c}_{j}^{k,s}$ are then projected to the original mask coordinate as $\bm{c}_{j}$ using Equations (17) and (18). Finally, the 2D coordinates are transformed into 3D space under the depth information $I_{\text{depth}}\in\mathbb{R}^{H\times W}$, thereby forming the initial constraint $\mathcal{C}_{i}^{\text{init}}$.
@@
where $\mathcal{N}_{4}\left(x,y\right)$ denotes the 4-neighborhood of pixel coordinate $(x,y)$, defined as
@@ III-C Real-Time Close-Loop Control Strategy
To solve the constrained optimization problem in Equation (2) with real-time performance, the Model Predictive Path Integral (MPPI) control algorithm is employed on the Isaac Gym simulation platform. MPPI is a stochastic optimization method designed for robot control in dynamic and uncertain environments. Essentially, the algorithm samples multiple candidate control sequences from the control input space, evaluates the cost functions based on simulated state trajectories, and updates the control policy through importance-weighted averaging. The GPU-accelerated parallel simulation enabled by Isaac Gym significantly enhances sampling efficiency, thereby supporting real-time execution in dynamic environments. For a detailed explanation of MPPI, please refer to [ref44, ref45]. In this work, we derive the specific formulation for solving the constrained optimization problem presented in Equation (2).
@@
where $\lambda_{p}$ and $\lambda_{r}$ are the weights for position and orientation costs, $\lambda_{c}$ denote the weight for the collision cost $\mathcal{D}_{c}$, which is formulated based on the ESDF representation, as described in [ref9]. In this work, the objective function is subject to preconditions and postconditions as specified in Equation (2), the precondition typically ensures that the end-effector is within the designated region prior to the execution stage, while the postcondition assesses whether the end-effector has successfully reached the desired pose, defined as
@@ III-D MLLM-driven Automated Modeling of TAMP
As depicted in Fig. 4, the Language Model Program (LMP) framework is built on a hierarchical recursive architecture [ref14], enabling structured invocation and multi-level prompt nesting. Based on the framework, we have designed and encapsulated seven LMP modules within ReSem3D, corresponding respectively to Task Planner, Constraint Extraction, Constraint Refinement, Subtask Execution, Preconditions Building, Cost Function Building, and PostConditions Building. Each module implements an independent prompt structure and interfaces with system components through standardized APIs, enabling the automated modeling for task and motion planning.
@@ IV-A Experimental Setup
Hardware Configuration. As shown in Fig. 6, our real-world experimental platform is constructed around the 6-DOF Universal Robots UR5e collaborative robotic arm, equipped with a PGI-140-80 two-finger parallel gripper and an Intel RealSense D435i RGB-D camera. For simulation, we employ the Omnigibson [ref46] platform developed by Stanford University, and utilize both the UR5e and the 7-DOF Franka Emika Panda collaborative arms.
@@
Baselines. We compare against three representative frameworks: VoxPoser [ref10], ReKep [ref9] and CoPa [ref2]. VoxPoser integrates large language and vision models to generate 3D value maps interpretable by robots, enabling closed-loop trajectory synthesis without predefined motion primitives. It leverages GPT-4 and a visual frontend for semantic grounding of task objects, supporting flexible manipulation across diverse tasks. ReKep fuses VLMs and MLLMs to infer semantic keypoint constraints for structured tasks and solves end-effector trajectories in real time via hierarchical optimization. CoPa leverages foundation vision-language models to propose a two-stage task execution process, performing task-oriented grasping via coarse-to-fine visual grounding in the first stage, and identifying task-relevant spatial and geometric constraints for robot planning in the second stage. All three frameworks deeply integrate large language and vision models, employing semantics-driven visual grounding and control optimization to achieve efficient and robust robotic manipulation.
@@
VFM and MLLM. We utilize GPT-4o [4] from OpenAI as a callable MLLM and combine it with the visual foundation model FastSAM [ref5] to enable high-efficiency multimodal reasoning (prompt design is available at ReSem3D.github.io. During the dynamic disturbance phase, we incorporate the TAPNext model from Google DeepMind [ref47] together with real-time RGB frame input to achieve online point tracking at approximately 20 Hz. This facilitates real-time inference of evolving spatial constraints in dynamic environments, thereby enhancing the stability and robustness of task execution.
IV-B Evaluation of ReSem3D in Simulated Environments